can you help me name the property illstrated by each statement
5. 6 (u + 2v) = 6u + 12v

2. 5(m)(n) =5(m)n
7. (2a + 3b) + 4c = 2a + (3b +4c)

4. 3t + 2r = 2r + 3t
6. 0 = 100(0)
9. gx =xg
8. pq + n = qp + n
10. 15(c + d) = 15(d + c )

Answer by chitra(359)

here are the solutions..

6 (u + 2v) = 6u + 12v ==> distributive property


5(m)(n) =5(m)n ==> associative property


(2a + 3b) + 4c = 2a + (3b +4c) ==> Associative property


3t + 2r = 2r + 3t ==> Commutative property for addition.

0 = 100(0) ==> multiplication by zero

gx = xg ==> Commutative Property

pq + n = qp + n Commutative

15(c + d) = 15(d + c ) Commutative prop of Addition.


Hence, the solution.

regs